OpenStudy (anonymous):

A sample contains 6.444 grams of boron and 1.803 grams of hydrogen what is the empirical of the compound?

OpenStudy (anonymous):

moles of B in the sample = 6.444 / 10.8117 = 0.596 moles moles of H = 1.803 / 1.0079 = molar ratio of B : H = 0.596:1.7888 or 1:3 empirical formula is BH3